OpenStudy (anonymous):

Which of the following is the equation of the vertical line that passes through. (4,6) A. Y=6 B. y=4 C. X=6 D. X=4

OpenStudy (anonymous):

Equations of vertical lines are always described using x. Since the point is (4,6) , the equation of the vertical line passing through that point should be x=4.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problem! :)

OpenStudy (anonymous):

Can u help me with this one? Write an equation in y=mx+b form for the table # of books purchased -1-2-3-4 Cost of book club. -21-30-39-48 A. Y=3x+21 B. y=9x+12 C. Y=5x+21 D. Y=7x+12

OpenStudy (anonymous):

@jkp269

OpenStudy (anonymous):

Let me see…

OpenStudy (anonymous):

I'm thinking that the # of books purchased are the x-values, and "Cost of book club" contains the y-values.

OpenStudy (anonymous):

We can get two points from this table which can help us to find the slope.

OpenStudy (anonymous):

So I'll choose (1,21) and (2,30) The formula to find the equation using two points is: y-y1=m(x-x1) m=(y2-y1)/(x2-x1) Label your points so that it will be easy to plug them in. x1=1 y1=21 x2=2 y2=30 So m=(30-21)/(2-1)=9/1=9 Now choose one point and plug into the formula: y-21=9(x-1) Simplify y=9x-9+21 y=9x+12

OpenStudy (anonymous):

Thanks again.